# Conformal four-point ladder integrals in diverse dimensions and polylogarithms

### v1

@@ Conclusions
Acknowledgments. We are grateful to A.C. Petkou for drawing our attention to the problem of evaluating conformal ladder diagrams in diverse dimensions. We would also like to thank A.I. Davydychev, A.V. Kotikov and A.C. Petkou, for useful comments and stimulating discussions. The work of S.E.D. and A.P.I. was supported by the grant RSF No. 23-11-00311.
@@ Appendix A The limit $\lambda\to 0$ for the function $\widetilde{\Phi}^{(\beta)}_{L}(z,\bar{z};\lambda)$
Thus, taking into account the generating function (2.37) of the Gegenbauer polynomials $C^{(\lambda)}_{n}(r)$, we find the expansion of $C^{(\lambda)}_{n}(r)$ for $\lambda\to 0$:
@@
It is noteworthy that the expression (A.5) can be written in a more compact form by means of the symmetry property of the eigenvalue (2.11) for $\lambda=0$:
@@
Then, we perform the change of variables $n\mapsto-n$ in the right-hand side of (A.5) in the sum containing $\bar{z}^{n}$, use property (A.6) and finally obtain concise formula
@@
can be easily checked. Indeed, using reflection formula $\Gamma(x)\Gamma(1-x)=\pi/\sin(\pi x)$ for the $\Gamma$-function, one can show that
@@ Appendix B Explicit check of the loop shift identity for $\lambda=1,\beta=1$
This appendix is dedicated to explicit check of the identity (2.54). We use the representation (3.26) for the function $\tilde{\Phi}^{(1)}_{L}(z,\bar{z};1)$ via classical polylogarithms
@@
Applying operator $\left(R^{(1)}_{\ell}(1)\right)^{-1}$ in the form (2.53) with $\beta=1$ we have
@@
where we used the following formula
@@ Appendix C Symmetrical polynomials in $z,\bar{z}$ in the case $\beta=1$
In this appendix, we discuss the properties of the rational functions arising as a coefficients of classical polylogarithms in the answer for ladder diagrams with arbitrary positive integer $\lambda$ and $\beta=1$ (this corresponds to the ladder diagrams Fig. 1 in $D=2\lambda+2$ dimensions with indices $1$ on the horizontal lines and D/21D/2-1 on the vertical lines). In what follows we summarize and generalize properties noted in the examples in section 3.1.
@@
where polynomial $G_{k}^{(\lambda)}(z,\bar{z})$ can be written in the form
@@
Proof. We start with the relation (3.17) and use representation in the r.h.s. of (3.16) to express the operator $P_{\lambda}(z\partial_{z})$ (below we use the notation from (3.22))
@@
In order to use the property (3.23) to express the derivative of polylog we rewrite the derivative in the form
@@
which follows from the fact $z^{-k}(z\partial_{z}-k+1)=\partial_{z}z^{-k+1}$. Introducing auxiliary variable $\alpha$ we rewrite
@@
Putting this relation in (C.3) and expanding $\alpha^{z\partial_{z}}$ in Taylor series we get
@@
Now we use (3.23) and calculate sum over $n$ which results in
@@
Note that we introduce an auxiliary variable $x$ to underline that one should treat operators in bracket as commuting and put $x=z$ after applying all derivatives. Also note that if $p>\lambda-1$ the corresponding terms in the sum nullifies after substituting $\alpha=1$. Expression (C.6) allows us to single out the contribution proportional to the polylogs and get the following representation for the polynomial (C.2)
@@
Rescaling auxiliary variable $t\mapsto tz$ we conclude (C.2). ∎
@@
Function $G_{k}^{(\lambda)}(z,\bar{z})$ defined in (C.2) is a homogeneous polynomial of degree $\lambda-1-k$ symmetrical under the transformation $z\leftrightarrow\bar{z}$ if $\lambda\in\mathbb{Z}_{>0}$ and $k=0,\ldots,\lambda-1$.
Proof. We organize the proof of this proposition in three steps. First, we show that $G_{k}^{(\lambda)}(z,\bar{z})$ is a polynomial. Indeed, application of the derivative with repsect to $t$ can be rewritten as
@@
Note that first term
$\partial_{t}^{n}\Big[(1+t)^{\lambda-1}\log^{k}(1+t)\Big]\Big|_{t=0}=0,\hskip 17.00024pt\text{if}\penalty 10000\ \penalty 10000\ n\leq k.$
Thus, the highest power of $(z-\bar{z})$ in the denominator is $\lambda-1-k+\lambda=2\lambda-1-k$, which precisely cancels by the prefactor. Second, we show that polynomial $G_{k}^{(\lambda)}(z,\bar{z})$ is homogeneous. Introducing arbitrary parameter $\mu\in\mathbb{R}$ we conclude
@@
As the last step we show that $G_{k}^{(\lambda)}(z,\bar{z})=G_{k}^{(\lambda)}(\bar{z},z)$. In order to address the derivative at $t=0$ we rewrite (C.2) as a Cauchy integral
@@
where $\gamma$ is the infinitesimally small contour around $t=0$. Now we do the change of variables
@@
Also note that the contour of integration $\gamma$ maps onto itself. Applying change of variables (C.8) to the (C.7) we get
@@
Changing the order of terms in bracket we arrive at
@@ Appendix D Derivation of factorization in two-dimensional case
This appendix is dedicated to the calculation of integral over $\nu$ in (3.31) and further factorization in $z$ and $\bar{z}$. In principal, one can find a detailed derivation in [27, Section 5] but for convenience we repeat it here. The first step is the calculation of the integral over $\nu$ by residues which results in the following expression
@@
Let us comment on the calculation of residues. Assuming the closing of contour in the lower half-plane we need to calculate the residues of the function which contains poles of order $L+1$ at the points $\nu_{k}=-{\sf i}\frac{n+1}{2}-{\sf i}k,k=0,1,\ldots$ (see Remark 2 in section 2.1 and [27, Fig. 10]). The corresponding residue can be expressed as
@@
  • Use the reflection relations for gamma-function
@@
Next step is factorization. Using the evident change of summation index $p=n+k$ in the first sum we obtain
@@
Note that in the second sum inside brackets the factor $\Gamma^{-L-1}(1-p-\varepsilon)$ creates additional $\varepsilon^{L+1}$ so that the whole sum is annihilated by the operator $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$ and after all one obtains
@@ Appendix E Derivation of the loop recursion for the operator $\widetilde{R}_{\ell}$ in the case $\beta=1$
First we introduce instead of (2.31), (2.32) (for $\beta=1$) the function
@@
which also depends only on the conformal variables $z,\bar{z}$. The operator $(z\partial_{z}+\bar{z}\partial_{\bar{z}})$ commutes with $\dfrac{z+\bar{z}}{2\sqrt{z\bar{z}}}$ and its action to the function $\widetilde{\widetilde{\Phi}}_{L}^{(1)}(z,\bar{z})$ gives
@@
that proves the statement that operator (2.55) produces the recursion

### v2

@@ Acknowledgments
We are grateful to A.C. Petkou for drawing our attention to the problem of evaluating conformal ladder diagrams in diverse dimensions. We would also like to thank A.I. Davydychev, A.V. Kotikov and A.C. Petkou for useful comments and stimulating discussions.
@@ Appendix A The limit $\lambda\to 0$ for the function $\widetilde{\Phi}^{(\beta)}_{L}(z,\bar{z};\lambda)$
Thus, taking into account the generating function (2.37) of the Gegenbauer polynomials $C^{(\lambda)}_{n}(r)$, we find the expansion of $C^{(\lambda)}_{n}(r)$ as $\lambda\to 0$:
@@
It is noteworthy that expression (A.5) can be written in a more compact form by means of the symmetry property of the eigenvalue (2.11) for $\lambda=0$:
@@
Then, we perform the change of variables $n\mapsto-n$ in the right-hand side of (A.5) in the sum containing $\bar{z}^{n}$, use property (A.6), and finally obtain the concise formula
@@
can be easily checked. Indeed, using the reflection formula $\Gamma(x)\Gamma(1-x)=\pi/\sin(\pi x)$ for the $\Gamma$-function, one can show that
@@ Appendix B Explicit check of the loop shift identity for $\lambda=1,\beta=1$
This appendix is devoted to the explicit check of identity (2.54). We use the representation (3.26) for the function $\tilde{\Phi}^{(1)}_{L}(z,\bar{z};1)$ in terms of classical polylogarithms
@@
Applying the operator $\left(R^{(1)}_{\ell}(1)\right)^{-1}$ in the form (2.53) with $\beta=1$ we have
@@
where we used the following formula:
@@ Appendix C Symmetrical polynomials in $z,\bar{z}$ in the case $\beta=1$
In this appendix, we discuss the properties of the rational functions arising as coefficients of classical polylogarithms in the answer for ladder diagrams with arbitrary positive integer $\lambda$ and $\beta=1$ (this corresponds to the ladder diagrams Fig. 1 in $D=2\lambda+2$ dimensions with indices $1$ on the horizontal lines and D/21D/2-1 on the vertical lines). In what follows we summarize and generalize the properties noted in the examples in section 3.1.
@@
where the polynomial $G_{k}^{(\lambda)}(z,\bar{z})$ can be written in the form
@@
Proof. We start the relation (3.17) and use the representation in the r.h.s. of (3.16) to express the operator $P_{\lambda}(z\partial_{z})$ (below we use the notation from (3.22))
@@
In order to use the property (3.23) to express the derivative of polylog, we rewrite the derivative in the form
@@
which follows from the fact $z^{-k}(z\partial_{z}-k+1)=\partial_{z}z^{-k+1}$. Introducing an auxiliary variable $\alpha$ we rewrite
@@
Putting this relation in (C.3) and expanding $\alpha^{z\partial_{z}}$ in Taylor series, we get
@@
Now we use (3.23) and calculate the sum over $n$ which results in
@@
Note that we introduce an auxiliary variable $x$ to underline that one should treat operators in the brackets as commuting and put $x=z$ after applying all derivatives. Also note that if $p>\lambda-1$, the corresponding terms in the sum are nullified after substituting $\alpha=1$. Expression (C.6) allows us to single out the contribution proportional to the polylogs and get the following representation for the polynomial (C.2):
@@
Rescaling the auxiliary variable $t\mapsto tz$, we conclude (C.2). ∎
@@
THe function $G_{k}^{(\lambda)}(z,\bar{z})$ defined in (C.2) is a homogeneous polynomial of degree $\lambda-1-k$ symmetrical under the transformation $z\leftrightarrow\bar{z}$ if $\lambda\in\mathbb{Z}_{>0}$ and $k=0,\ldots,\lambda-1$.
Proof. We organize the proof of this proposition in three steps. First, we show that $G_{k}^{(\lambda)}(z,\bar{z})$ is a polynomial. Indeed, application of the derivative with respect to $t$ can be rewritten as
@@
Note that the first term
$\partial_{t}^{n}\Big[(1+t)^{\lambda-1}\log^{k}(1+t)\Big]\Big|_{t=0}=0,\qquad\text{if}~~n\leq k.$
Thus, the highest power of $(z-\bar{z})$ in the denominator is $\lambda-1-k+\lambda=2\lambda-1-k$, which is precisely canceled by the prefactor. Second, we show that the polynomial $G_{k}^{(\lambda)}(z,\bar{z})$ is homogeneous. Introducing an arbitrary parameter $\mu\in\mathbb{R}$, we conclude
@@
As the last step we show that $G_{k}^{(\lambda)}(z,\bar{z})=G_{k}^{(\lambda)}(\bar{z},z)$. In order to address the derivative at $t=0$, we rewrite (C.2) as a Cauchy integral
@@
where $\gamma$ is the infinitesimally small contour around $t=0$. Now we change the variables
@@
Also note that the contour of integration $\gamma$ maps onto itself. Applying the change of variables (C.8) to  (C.7), we get
@@
Changing the order of terms in the brackets, we arrive at
@@ Appendix D Derivation of factorization in two-dimensional case
This appendix is devoted to the calculation of integral over $\nu$ in (3.31) and further factorization in $z$ and $\bar{z}$. In principal, one can find a detailed derivation in (Derkachov et al., 2019, Section 5), but for convenience we repeat it here. The first step is the calculation of the integral over $\nu$ by evaluating residues, which results in the following expression
@@
Let us comment on the calculation of residues. Assuming the closing of contour in the lower half-plane, we need to calculate the residues of the function that contains poles of order $L+1$ at the points $\nu_{k}=-{\sf i}\frac{n+1}{2}-{\sf i}k,k=0,1,\ldots$ (see Remark 2 in section 2.1 and (Derkachov et al., 2019, Fig. 10)). The corresponding residue can be expressed as
@@
  • Use the reflection relations for the gamma-function
@@
The Next step is factorization. Using the evident change of the summation index $p=n+k$ in the first sum, we obtain
@@
Note that in the second sum inside the brackets the factor $\Gamma^{-L-1}(1-p-\varepsilon)$ creates additional $\varepsilon^{L+1}$ so that the whole sum is annihilated by the operator $\left.\partial_{\varepsilon}^{L}\right|_{\varepsilon=0}$ and after all one obtains
@@ Appendix E Derivation of the loop recursion for the operator $\widetilde{R}_{\ell}$ in the case $\beta=1$
First, we introduce instead of (2.31), (2.32) (for $\beta=1$) the function
@@
which also depends only on the conformal variables $z,\bar{z}$. The operator $(z\partial_{z}+\bar{z}\partial_{\bar{z}})$ commutes with $\dfrac{z+\bar{z}}{2\sqrt{z\bar{z}}}$ and its action on the function $\widetilde{\widetilde{\Phi}}_{L}^{(1)}(z,\bar{z})$ gives
@@
which proves the statement that operator (2.55) produces the recursion